A man of mass 60 kg record his weight on weighing machine placed inside lift The ratio of down with uniform speed of 4m/s will be weights of man recorded when lift is ascending up with uniform speed of 2 m/s to when it is descending down with uniform speed of 4m/s will be:--
  1. 0.5     b. 1      c. 2     d. none

Grade:11

7 Answers

Neeti
8 years ago
The language of the question is a little unclear. Could you post the question again in the comment box?
Ananya Sharma
8 years ago
A man of mass 60 kg record his wt. on a weigheing machine placed inside a lift. The ratio of weights of man recorded when lift is ascending up with uniform speed of 2m/s to when it is descending down with uniform speed of 4m/s will be:_
  1. 0.5      b. 1      c. 2      d. none
I want answer with full solution
Neeti
8 years ago
i think it’s option b because the lift is not undergoing accelaration so there is no pseudo force or accelaration therefore weight in both cases is the same.
Neeti
8 years ago
Neeti
8 years ago
 
I’m sorry i don’t know why the answer didn’t get posted. let me be more clear, The lift which is the frame of reference is not undergoing any accelaration therefore there is no pseudo force or pseudo accelration on the man therefore his original weight does not get altered. that is, in both cases weight will be 60N
Feel free to ask any follow up question and approve the answer if you like it :)
Arijeet Dey
6 years ago
Since the speed is uniform so there is not pseudo force in the lift so that the ratio is 1 unless the speed would be change
Satvik
6 years ago
When body goes up R1=m(g+a) R1=60(10+0) R1=600When body goes down R2=m(g-a) R2=60(10-0) R2=600 Ratio=R1/R2=600/600=1